# Support Recovery and $\ell_{2}$-Error Bound for Sparse Regression with Quadratic Measurements via Weakly-Convex-Concave Regularization

### v1

@@ 3 Optimization Algorithm
In this section, we discuss the numerical computation of the problem to assess the performance of the proposed method. Since $n$ and $d$ are given, to simplify notation we denote $\lambda_{n}=\lambda$, so that (2) is written as
@@
where $L(\bm{\beta}):=\frac{1}{4n}\sum_{i=1}^{n}(\bm{\beta}^{\top}\bm{Z}_{i}\bm{\beta}-y_{i})^{2}$ and $\lambda>0$. Noting the term $L(\bm{\beta})$ is smooth and the term $P_{\lambda}(\bm{\beta})$ is weakly convex-concave, we employ the PGA to solve the problem. The PGA is an established classical algorithm (Beck 2017) and has been applied in various inverse and optimization problems (e.g., Bolte et al. (2018), Soltanolkotabi (2019),Zhang et al. (2023), Fan et al. (2025b)). The core component of PGA is the proximal operator. Given a function $P:\mathbb{R}^{d}\to(-\infty,\infty)$, recall that the proximal operator associated with $P$ is defined as
@@
There exists a constant $r>0$, for any $\tau\in(0,\hat{L}^{-1}]$, with $\hat{L}=\sup_{\|\bm{\beta}\|\in B_{r}}\|\nabla^{2}L(\bm{\beta})\|$ and $B_{r}:=\left\{\bm{\beta}\in\mathbb{R}^{d}:\|\bm{\beta}\|\leq r\right\}$, such that for any minimizer $\hat{\bm{\beta}}$ of problem (2), it satisfies
@@
where the $j$th element of the vector $\hat{\bm{w}}_{\lambda}$ is defined as
@@
Here $\epsilon_{1}$ is any positive number.
These two characterizations provide the foundation for our algorithmic design. According to the relationship (9), the iterative scheme for solving (2) is outlined in Algorithm 1. For (10), the weight $\hat{\bm{w}}_{\lambda}$ depends on the unknown optimal solution $\hat{\bm{\beta}}$ rendering the problem circular and not directly solvable. To overcome this, we adopt the classic framework of iteratively reweighted $\ell_{1}$ algorithms (Zou (2006),Bai et al. 2024). Within this framework, we employ the Majorization-Minimization (MM) technique. At each iteration, the regularization term is linearly approximated via its first-order expansion, which constructs a surrogate function. This leads to a sequence of subproblems of the form
@@
It is worth noting that computing the step size $\tau_{k}$ is a crucial step, and its selection depends on the parameter $\hat{L}$, which is difficult to determine. To address this, we employ the Armijo line search method (11). The existence of such a smallest $j_{k}$ and the admissible range for $\tau_{k}$ are established in Lemma 5 of Appendix B.
Having established the step size selection strategy, we now turn to the convergence analysis of the algorithms.
@@
Assume that $\left\{\bm{\beta}^{k}\right\}$ is the sequence generated by Algorithm 1 or 2. Then the following conclusions hold,
@@
    Every accumulation point of the sequence $\{\bm{\beta}^{k}\}$ generated by Algorithm 1 satisfies the fixed-point equation (9), and that generated by Algorithm 2 satisfies (10).
@@ 4 Numerical Experiments
In this section, we examine the finite sample performance of the proposed method through numerical simulations. All experiments are implemented in MATLAB(R2023b) on a laptop with 32GB memory. The MATLAB code is available at https://github.com/fjmath/sparseQMR_WCCP.
We compares the performance of the SCAD, MCP, Firm, LOG, and EXP regularizers with the $\ell_{1}$ and $\ell_{1/2}$ regularizers. The initial point is obtained using the sparse spectral initialization method proposed by Chen et al. (2025). For all regularization methods, the optimal step size is determined using the Armijo line search. The regularization parameter is chosen according to the scheme of Chen and Ng (2022), defined as $\lambda_{n}=\sqrt{\frac{c\ln(d)}{n^{2}}\sum_{i=1}^{n}(\bm{\beta}^{\top}\bm{Z}_{i}\bm{\beta}-y_{i})^{2}}\|\bm{\beta}\|$, where the coefficient $c$ is determined via cross-validation. The relative error in the experiments is computed as follows,
@@
Experiment 1. The sample size is $n$, and the measurement matrix $\bm{Z}$ is taken to be a symmetric matrix with Gaussian entries. The noise $\bm{\varepsilon}$ follows a sub-Gaussian distribution. The true signal $\bm{\beta}^{\star}$ is generated as a Gaussian random vector. This experiment compares the performance of WCCP-regularized with the $\ell_{1}$ and $\ell_{1/2}$ regularizers under varying $n/d$ ratios. In the experiments, noise $\varepsilon_{i}\sim\sigma\mathcal{N}(0,1)$ is added to the measurements $y_{i}$. Regarding the signal dimension $n$ and sparsity level $s$, we consider two cases: d=128,s=10d=128,~s=10 and d=256,s=15d=256,~s=15. For $\sigma=0.01,~\text{and}~0.05$, the recovery of the unknown signal $\bm{\beta}^{\star}$ is considered successful when the error is below $10^{-3}~\text{and}~10^{-2}$ respectively. The measurement ratios are set as $n/d=\left\{0.1,\dots,0.9,1\right\}$, with each experiment run 100 times. The performance is evaluated based on success rate and average relative error, with the latter is calculated only for successful trials. Furthermore, Supplementary Tables S1–S8 confirm that as $n/d$ approaches 1, several WCCP regularizers—particularly SCAD and MCP—achieve lower relative errors and higher stability than $\ell_{1}$ and $\ell_{1/2}$, while being orders of magnitude faster. In contrast, Firm, LOG and EXP incur substantially higher computational costs. Overall, SCAD and MCP offer the best trade‑off between accuracy and efficiency.
Figure 1 illustrates the curves of success rate as functions of the ratio $n/d$. The results confirm that WCCP methods generally outperform $\ell_{1}$ regularization and are competitive with $\ell_{1/2}$. The SCAD and MCP penalties achieve the highest success rates, especially as $n/d$ increases. Other WCCP members (Firm, LOG, and EXP) also exhibit strong performance, with LOG and EXP proving particularly robust across noise levels. Detailed mean squared error and variance results for all measurement ratios are provided in Appendix C (Tables 5-12). These results show that as $n/d$ approaches 1, our works exhibits superior relative-error performance. Therefore, our method is more stable in identifying the sparse support.
@@
Experiment 2. This experiment demonstrates the performance of WCCP-regularized method in the sparse phase retrieval problem. The proximal operator of the EXP regularizer is computationally expensive; therefore, Algorithm 2 is used to solve the problem in this experiment. The sparse test signal is generated from the standard test image "cameraman.tif" (available in MATLAB’s image processing toolbox) following the wavelet-based sparsification procedure Jagatap and Hegde (2018) and Shah and Hegde (2021). Our implementation adapts the publicly available code from Shah and Hegde (2021): we resize the image to 64×64 pixels, apply a 4-level Haar wavelet transform, and retain only the top $5\%$ of the coefficients by magnitude. This differs from Shah and Hegde (2021) in both the test image used and the sparsity criterion (percentage-based vs. fixed count). In the experiments, noise $\varepsilon_{i}\sim\sigma\mathcal{N}(0,1)$ is added to the measurements.
The comprehensive numerical experiments demonstrate the clear advantages of WCCP-regularized methods over traditional approaches for sparse phase retrieval. As shown in Tables 1 and 3, WCCP methods, particularly SCAD and MCP, consistently achieve near-perfect reconstruction (SSIM $\approx$ 0.9992) across all noise levels and measurement ratios, while traditional $\ell_{1}$ and $\ell_{0.5}$ methods show significant performance degradation, especially under high noise conditions and low measurement ratios. Notably, MCP achieves perfect reconstruction at measurement ratio $n/d=0.4$, whereas $\ell_{1}$ requires $n/d\geq 0.9$ to reach comparable quality, demonstrating the superior sparse recovery capability of WCCP regularizers.
In addition to enhanced reconstruction quality, WCCP methods offer substantial computational advantages. As evidenced in Tables 2 and 4, SCAD and MCP complete reconstruction tasks in significantly less time compared to $\ell_{1}$ regularization, while maintaining superior accuracy. For instance, at noise level $\sigma=10^{-4}$, $\ell_{1}$ requires 139.08 seconds while SCAD and MCP complete in only 0.61 and 0.58 seconds respectively. These results confirm that WCCP-regularized methods provide an efficient and robust framework for sparse signal recovery, making them particularly suitable for practical applications requiring high-quality reconstruction under challenging conditions.
@@
Table 1: Structural Similarity Index (SSIM) performance of sparse regularization methods at different noise levels (measurement ratio $n/d=0.50$)
@@ 0 $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ 1
0 $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ 1
L1 1.7725 1.9901 139.0842 1.9897 1.8820 1.9666 1.3740
SCAD 0.5458 0.6096 0.6105 0.5703 0.5774 0.5958 0.7209
MCP 0.5206 0.5516 0.5760 0.5341 0.5313 0.7673 0.6821
$L_{0.5}$ 0.4745 0.5401 0.5185 0.5150 0.4807 0.4364 0.3899
Firm 0.6950 0.7791 0.7060 0.6844 0.6630 84.0347 0.5769
LOG 1.1945 1.3752 1.3729 1.1638 1.2105 1.3171 0.9131
EXP 1.1392 15.2622 1.1028 1.1917 1.1054 1.2542 0.8009
Table 3: Structural Similarity Index (SSIM) performance at different measurement ratios (noise level $\sigma=10^{-2}$)
@@
0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
L1 0.8037 1.5078 1.9668 1.9321 2.0649 2.0567 2.1940 2.3576 2.3357
SCAD 0.3875 0.4539 0.5693 0.5984 0.6111 0.6245 0.6540 0.6864 0.7105
MCP 0.3673 0.4219 0.5328 0.5630 0.5906 0.5945 0.6095 0.6428 0.6673
$L_{0.5}$ 0.3269 0.3634 0.4327 0.4728 0.5130 0.5303 0.5673 0.5848 0.6059
Firm 0.4306 0.5050 0.6558 0.6647 0.6916 0.7159 0.7635 0.7762 0.8061
LOG 0.7325 0.8526 1.0285 1.1950 1.3259 1.4074 1.5055 1.5480 1.6134
EXP 0.7063 0.8373 1.0086 1.1550 1.2877 1.3680 1.4670 1.5155 1.5763

### v2

@@ 3 Optimization Algorithm
In this section, we discuss the numerical computation of the problem to assess the performance of the proposed method. Since $n$ and $d$ are fixed, we drop them in $\lambda_{n}=\lambda$ and write (2) as
@@
where $L(\bm{\beta}):=\frac{1}{4n}\sum_{i=1}^{n}(\bm{\beta}^{\top}\bm{Z}_{i}\bm{\beta}-y_{i})^{2}$ and $\lambda>0$. Since $L(\bm{\beta})$ is smooth and $P_{\lambda}(\bm{\beta})$ is weakly convex-concave, we employ the PGA to solve the problem (8). The PGA is an established classical algorithm (Beck 2017) and has been applied in various inverse and optimization problems (e.g., Bolte et al. (2018), Soltanolkotabi (2019),Zhang et al. (2023), Fan et al. (2025b)). The core component of PGA is the proximal operator. Recall that, given a function $P:\mathbb{R}^{d}\to(-\infty,\infty)$, the proximal operator associated with $P$ is defined as
@@
There exists a constant $r>0$, such that for any $\tau\in(0,\hat{L}^{-1}]$ and minimizer $\hat{\bm{\beta}}$ of problem (2), it holds
@@
where $\hat{L}=\sup_{\|\bm{\beta}\|\in B_{r}}\|\nabla^{2}L(\bm{\beta})\|$, $B_{r}:=\left\{\bm{\beta}\in\mathbb{R}^{d}:\|\bm{\beta}\|\leq r\right\}$ and $\hat{\bm{w}}_{\lambda}$ is a vector with $j$th element
@@
Here $\epsilon_{1}$ is a positive number.
The above two characterizations provide the foundation for our algorithmic design. While the iterative scheme for solving (2) is outlined in Algorithm 1 according to (9), the weight $\hat{\bm{w}}_{\lambda}$ in (10) depends on the optimal solution $\hat{\bm{\beta}}$ which is unknown. To overcome this issue, we adopt the framework of iteratively reweighted $\ell_{1}$ algorithms (Zou (2006),Bai et al. 2024). Within this framework, we employ the Majorization-Minimization (MM) technique, where, at each iteration, the regularization term is linearly approximated via its first-order expansion, which constructs a surrogate function. This leads to a sequence of sub-problems of the form
@@
Note that computing the step size $\tau_{k}$ is crucial, and its selection depends on the parameter $\hat{L}$ that is difficult to determine. To address this issue, we employ the Armijo line search method (11). The existence of a smallest $j_{k}$ and the admissible range for $\tau_{k}$ are established in Lemma 5 of Appendix B. Having established the step size selection strategy, we now turn to the convergence analysis of the algorithms.
@@
Let $\left\{\bm{\beta}^{k}\right\}$ be the sequence generated by Algorithm 1 or 2.
@@
    Every accumulation point of $\{\bm{\beta}^{k}\}$ generated by Algorithm 1 satisfies the fixed-point equation (9), and that generated by Algorithm 2 satisfies (10).
@@ 4 Numerical Experiments
In this section, we examine the finite sample performance of the proposed method through numerical simulations. All experiments are implemented using MATLAB(R2023b) and the code is available at https://github.com/fjmath/sparseQMR_WCCP.
We compare the performance of several WCCP, including SCAD, MCP, TL1, LOG, and EXP regularizers, with $\ell_{1}$ and $\ell_{1/2}$ regularizers. To enable a fair comparison across penalties where $\mu$ scales differently with $\lambda_{n}$, we introduce the standardized concavity
$\kappa:=\mu/\lambda_{n},$
a dimensionless constant determined solely by each penalty’s shape parameters (Table 1). This quantity captures the intrinsic curvature of the penalty and aligns with the maximum concavity measure of Zhang (2010), expressed here within the $\mu$-amenable framework of Loh and Wainwright (2017).
Table 1: Shape parameters and standardized concavity $\kappa=\mu/\lambda_{n}$ for the seven penalties, ordered by increasing concavity.
Penalty $\ell_{1}$ MCP SCAD TL1 EXP LOG $\ell_{1/2}$
Shape param. $\gamma=3$ $a=3.7$ $a=3$ $\sigma=0.5$ $\varepsilon=0.1$ $q=0.5$
$\kappa$ 0 $\frac{1}{\gamma}$ $\frac{1}{a-1}$ $\frac{2(a+1)}{a^{2}}$ $\frac{1}{\sigma^{2}}$ $\frac{1}{\varepsilon^{2}}$ $\infty$
[…]
The initial point is obtained using the sparse spectral initialization method proposed by Chen et al. (2025). For all regularization methods, the optimal step size is determined using the Armijo line search. The regularization parameter is chosen according to the scheme of Chen and Ng (2022), defined as $\lambda_{n}=\sqrt{\frac{c\ln(d)}{n^{2}}\sum_{i=1}^{n}(\bm{\beta}^{\top}\bm{Z}_{i}\bm{\beta}-y_{i})^{2}}\|\bm{\beta}\|$, where the coefficient $c$ is determined via cross-validation. The relative error in the experiments is computed as follows,
@@
Performance is then evaluated by
  • Success rate: proportion of trials with $\text{RelErr}<10^{-3}$;
  • Support recovery: true positive rate (TPR), false positive rate (FPR), and F1 score;
  • Bias for large coefficients: average relative deviation $|(\hat{x}_{j}-x_{j})/x_{j}|$ over the large‑coefficient support;
  • Computational time.
4.1 Experiment 1: Synthetic quadratic measurements
We first study recovery performance using randomly generated quadratic measurements. The sensing matrices $\bm{Z}_{i}$ are generated as symmetric matrices with Gaussian entries, and the noise vector $\bm{\varepsilon}$ follows a sub-Gaussian distribution. The true signal $\bm{\beta}^{\star}$ is generated as a sparse Gaussian vector.
Additive noise $\varepsilon_{i}\sim\sigma\mathcal{N}(0,1)$ is added to the measurements $y_{i}$. Two problem settings are considered $(d,s)=(128,10)$ and $(d,s)=(256,15).$ The sampling ratios are chosen as $n/d\in\{0.1,0.2,\dots,1.0\},$ and each configuration is repeated 100 times. Recovery is regarded as successful when $\text{RelErr}<10^{-3}$.
Figure 1 reports the success rates as functions of the sampling ratio $n/d$. Across all experimental settings, WCCP substantially outperform the convex $\ell_{1}$ regularization in the moderate undersampling regime ($0.3\leq n/d\leq 0.6$). For example, when $n/d=0.4$ with $d=128$, $s=10$, and $\sigma=0.01$, the success rates of WCCP range from $0.30$ (MCP) to $0.72$ (TL1), while the $\ell_{1}$ method barely exceeds $0.01$.
Among the WCCP methods, TL1 consistently achieves the highest success rates and often reaches $0.9$ once $n/d\approx 0.5$ under low noise. LOG and EXP also exhibit strong performance and closely follow TL1. The $\ell_{1/2}$ penalty shows mixed behavior relative to $\ell_{1}$: although it sometimes attains slightly higher F1 scores due to lower false positive rates, its success rate and estimation accuracy remain inferior to those of the WCCP methods.
Detailed statistics (mean and standard deviation) for RelErr, TPR, FPR, F1 score, bias for large coefficients, estimated sparsity, and computational time are reported in Appendix C (Tables 6–21). The results indicate that WCCP methods consistently achieve lower estimation error, improved support recovery, and reduced shrinkage bias compared with $\ell_{1}$, particularly when the sampling ratio is moderate.
These results demonstrate that weakly convex concave penalties significantly improve sparse recovery under quadratic measurements, especially in undersampled regimes. In particular, TL1 and LOG provide the most favorable balance between recovery accuracy and sparsity control.
@@ 4.1 Experiment 1: Synthetic quadratic measurements
4.2 Experiment 2: Sparse phase retrieval
The second experiment evaluates the proposed approach on sparse phase retrieval. Because the proximal operator of the EXP regularizer is computationally expensive, Algorithm 2 is used to solve the corresponding optimization problem.
The sparse signal is generated from the standard test image cameraman.tif. Following the wavelet-based sparsification procedure of Shah and Hegde (2021), the image is resized to $64\times 64$ pixels and transformed using a four-level Haar wavelet transform. Only the largest $5\%$ of coefficients are retained to form the sparse representation. This differs from Shah and Hegde (2021), which retains a fixed number of coefficients.
Measurement noise is generated as $\varepsilon_{i}\sim\sigma\mathcal{N}(0,1)$.
Tables 2 and 4 report reconstruction quality measured by the structural similarity index (SSIM). Several WCCR penalties, including SCAD, MCP, TL1, and EXP, consistently achieve near-perfect reconstruction (SSIM $\approx 0.9992$) across a wide range of measurement ratios and noise levels. In contrast, the classical $\ell_{1}$ regularization shows noticeable degradation when the number of measurements is limited. For example, at $n/d=0.4$, MCP already attains nearly perfect reconstruction, while $\ell_{1}$ achieves only SSIM =0.708=0.708.
The $\ell_{1/2}$ penalty is computationally efficient but appears less stable when the number of measurements is small, where reconstruction quality exhibits larger variability.
Tables 3 and 5 summarize the computational time. SCAD and MCP achieve competitive or faster runtimes than $\ell_{1}$ while maintaining superior reconstruction accuracy. For instance, at $n/d=0.2$, SCAD and MCP require $0.62$ and $0.38$ seconds respectively, compared with $0.97$ seconds for $\ell_{1}$. Although TL1 also achieves near-perfect reconstruction, its computational cost is substantially higher.
These results confirm that the advantages of WCCP extend to practical imaging problems. In particular, SCAD and MCP achieve highly accurate reconstructions with competitive computational cost, whereas TL1 attains excellent reconstruction quality at the expense of increased computation.
@@ 4.2 Experiment 2: Sparse phase retrieval
Table 2: SSIM performance of sparse regularization methods at different noise levels (measurement ratio $n/d=0.50$)
@@ 0 $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
0 $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
$\ell_{1}$ 2.5091 (0.0971) 2.8332 (0.5230) 2.7009 (0.4100) 3.4049 (0.3444) 2.4327 (0.0431) 2.4057 (0.0767)
$\ell_{1/2}$ 0.7674 (0.1992) 0.9997 (0.5615) 1.2502 (0.8086) 1.1966 (0.7994) 0.7921 (0.3556) 0.8601 (0.5791)
SCAD 2.5128 (0.1457) 2.7934 (0.5774) 2.7956 (0.4786) 3.3247 (0.4467) 2.4365 (0.1589) 2.3920 (0.1716)
MCP 1.9420 (0.3407) 2.3035 (0.7515) 2.3677 (0.5875) 2.5085 (0.5659) 1.7920 (0.3211) 1.6350 (0.4330)
TL1 7.0248 (1.0057) 7.5769 (1.1488) 7.0298 (1.6241) 7.0558 (1.0862) 6.6681 (1.0291) 6.6241 (1.0589)
LOG 0.7942 (0.4062) 0.9953 (0.6267) 1.2440 (0.9794) 1.1928 (0.8394) 0.8852 (0.5659) 0.8874 (0.6563)
EXP 1.7747 (0.2575) 2.1511 (0.5463) 2.0821 (0.4885) 2.2267 (0.4384) 1.6534 (0.2419) 1.5930 (0.2972)
Table 4: SSIM performance at different measurement ratios (noise level $\sigma=10^{-2}$)
@@
0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
$\ell_{1}$ 0.5246 (0.0184) 0.9662 (0.0667) 1.4456 (0.0270) 2.2435 (0.4514) 3.4744 (0.1729) 4.0473 (0.1358) 4.6626 (0.1794) 4.5466 (1.5119) 4.5514 (0.4814) 4.3822 (0.5646)
$\ell_{1/2}$ 0.6842 (0.0559) 1.1321 (0.0987) 1.5254 (0.1543) 1.3441 (0.5833) 1.3820 (0.9612) 0.9934 (0.4020) 0.8612 (0.1362) 0.7285 (0.2167) 0.7256 (0.1145) 0.7511 (0.0697)
SCAD 0.2307 (0.0683) 0.6189 (0.1964) 1.1997 (0.3138) 2.3479 (0.5072) 3.4289 (0.3615) 3.8048 (0.4553) 3.7234 (0.6130) 2.9868 (0.7921) 2.9846 (0.6118) 2.9563 (0.7427)
MCP 0.1622 (0.0460) 0.3774 (0.1357) 0.8955 (0.3318) 2.1388 (0.5367) 2.6129 (0.5691) 2.7603 (0.7117) 2.7807 (0.8023) 2.1441 (0.6194) 2.1697 (0.4512) 2.2514 (0.5815)
TL1 2.7968 (0.5581) 2.1315 (0.6213) 4.2715 (1.8003) 7.3948 (0.3143) 7.3901 (1.0583) 7.2175 (1.5624) 6.0507 (1.6447) 7.5181 (6.8340) 5.4924 (0.5665) 5.1223 (0.8201)
LOG 0.5163 (0.0153) 0.9778 (0.0968) 1.3788 (0.1775) 1.7594 (0.7271) 1.1436 (0.6788) 1.0426 (0.7505) 0.7997 (0.1000) 0.7236 (0.2294) 0.6966 (0.0640) 0.7166 (0.0405)
EXP 0.5661 (0.0163) 0.9938 (0.0654) 1.4315 (0.0442) 2.1014 (0.4496) 2.4284 (0.4303) 2.4131 (0.4705) 2.4388 (0.8907) 1.9862 (0.7783) 1.9755 (0.2541) 1.9815 (0.2683)
4.3 Discussion.
To summarize, the above experiments consistently demonstrate the advantages of WCCP for sparse quadratic measurements model. Compared with classical $\ell_{1}$ and $\ell_{1/2}$ regularization, WCCP methods achieve more accurate support recovery, lower estimation error, and substantially reduced shrinkage bias, while maintaining competitive computational efficiency. These improvements are particularly evident in the moderately undersampled regime, where convex regularization tends to suffer from large bias and excessive false positives.
The empirical results also reveal a clear relationship between the concavity of the penalty functions and their recovery performance. As characterized by the standardized concavity parameter $\kappa$, penalties with stronger concavity generally provide a closer approximation to the $\ell_{0}$ regularizer and therefore exhibit improved support identification and reduced estimation bias. In particular, TL1 and LOG, which possess relatively stronger concavity among the considered penalties, consistently achieve the highest success rates and the most accurate support recovery across a wide range of sampling ratios. By contrast, the convex $\ell_{1}$ penalty corresponds to zero concavity and tends to introduce significant shrinkage bias, often leading to a large number of false positives. Penalties with moderate concavity, such as SCAD and MCP, offer a favorable compromise between statistical accuracy and computational efficiency, producing stable recovery performance while remaining relatively inexpensive computationally.
From a practical perspective, different penalties provide distinct trade-offs between recovery accuracy and computational cost. TL1 and LOG tend to perform best in strongly undersampled regimes where accurate support recovery is critical. SCAD and MCP, on the other hand, achieve competitive reconstruction quality with significantly lower computational cost and may therefore be preferable in large-scale applications. The $\ell_{1/2}$ penalty is computationally efficient but appears less stable when the number of measurements is small.
Overall, these empirical findings are consistent with the theoretical properties established earlier in the paper. In particular, the results support the theoretical insight that WCCP provides a closer approximation to the $\ell_{0}$ penalty while preserving tractable optimization. This combination of favorable statistical properties and practical computational behavior makes WCCP a promising framework for sparse quadratic measurements model.